# Sustainable Exploitation Equilibria for Dynamic Games

### v1

@@ Abstract
We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure. SEE imposes two additional discipline conditions: (i) viability, requiring state trajectories to remain inside a sustainability set; and (ii) renegotiation-proofness with exploiter-optimal selection, to retain only those viable equilibria that are immune to Pareto-improving renegotiations, with ties resolved in favor of the exploiter. In our base formulation the exploitee cannot exit the relationship (no outside option), but retains a strategic effort margin that affects dynamics and payoffs. We establish existence under appropriate conditions and illustrate SEE in a hegemon-client model of foreign politics, where tribute demands trade off against the client’s governance effort.
@@ 1 Introduction
Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself. Standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter extracts too aggressively, driving the state outside the range in which the exploitee can survive. To address this, the paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary MPE that selects only those equilibria that (i) keep the state within a viability set and (ii) are immune to Pareto-improving renegotiations, with exploiter-optimal selection. The contribution is to formalize a simple and tractable equilibrium notion suited to persistent asymmetric relationships in which collapse is feasible but avoided in equilibrium. SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics. In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee.
@@ 3 Model and Definition
Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Let $\mathcal{E}$ denote the set of all stationary Markov Perfect Equilibria, and let $\mathcal{E}_{V}\subseteq\mathcal{E}$ denote the subset whose induced state process remains in $V$. We assume $\mathcal{E}_{V}\neq\emptyset$.
@@ 4 Relation to Nash and Markov Perfect Equilibrium
5 Existence
Theorem 1 (Existence of SEE).
Suppose the state space $S$ and action sets $X(s),E(s)$ are compact; payoffs are bounded and continuous; and the transition kernel $Q$ satisfies the Feller property. Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Assume that there exists at least one stationary Markov Perfect Equilibrium whose induced state process remains in $V$. Then:
(i) the set of viable stationary Markov Perfect Equilibria is nonempty and compact;
(ii) the subset of renegotiation-proof viable equilibria is nonempty; and
(iii) an exploiter-optimal selection exists.
Hence a Sustainable Exploitation Equilibrium (SEE) exists.
@@ 5 Existence
(1) Existence of stationary MPE. Under compact state and action spaces, bounded and continuous payoffs, and Feller continuity of the transition kernel, the stationary best-response correspondence is nonempty, convex-valued, and upper hemicontinuous. By Glicksberg’s fixed point theorem, there exists at least one stationary Markov Perfect Equilibrium (Maskin and Tirole, 2001).
(2) Viability. Let $\mathcal{E}$ denote the set of stationary MPE and define $\mathcal{E}_{V}\subseteq\mathcal{E}$ as the subset whose induced state process remains in the invariant compact set $V$. By assumption, $\mathcal{E}_{V}$ is nonempty. Compactness of $S$ and continuity of payoffs and transitions imply that the associated set of equilibrium value functions for $\mathcal{E}_{V}$ is compact in the product topology.
(3) Renegotiation-proofness. The set of feasible continuation values generated by $\mathcal{E}_{V}$ is compact. Therefore there exists at least one equilibrium in $\mathcal{E}_{V}$ that is not strictly Pareto-dominated by any other viable equilibrium; such an equilibrium is renegotiation-proof.
(4) Selection. The exploiter’s value function is continuous on the nonempty compact set of renegotiation-proof viable equilibria and hence attains a maximum. Any maximizer is a Sustainable Exploitation Equilibrium. ∎
@@ 6.6 Stationary SEE and regimes
A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Two regimes arise:
@@
Additional Remark on Outside Option Expansions
One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$. This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit. We deliberately exclude outside options in the baseline SEE definition, since the core notion of exploitation is precisely that the exploitee cannot walk away. In other settings, the SEE framework can be adapted straightforwardly by adding the IR condition, producing a refinement that nests our baseline definition, but applies to a broader class of sustainable strategic relationships.
@@ 7 Conclusion
With a non-exiting but strategically responsive exploitee, SEE refines MPE by imposing viability and renegotiation-proof exploiter-optimal selection. In the hegemon–client example, SEE yields restrained extraction that keeps the regime just at the viability boundary. More generally, the framework applies whenever one party cannot exit but retains an effort margin, including colonial settings, labor relationships and locked-in suppliers. By combining dynamic sustainability with renegotiation-proofness, SEE offers a compact refinement of MPE for analyzing persistent exploitative interactions in economics and political economy.

### v2

@@ Abstract
We introduce the Sustainable Exploitation Equilibrium (SEE), a refinement of Markov Perfect Equilibrium (MPE) for dynamic games with an exploiter–exploitee structure. SEE imposes renegotiation-proof exploiter-optimal selection on the set of rationally viable stationary Markov equilibria, where viability follows from sequential rationality when exiting a sustainability set entails catastrophic losses. Unlike MPE, SEE rules out equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs. The exploitee cannot exit, but retains a strategic effort margin affecting dynamics and payoffs. We establish existence under standard conditions, and the refinement is illustrated in a hegemon–client model of foreign politics.
@@ 1 Introduction
Exploitation arises when one agent extracts value from another whose continued functioning is essential for the relationship itself. In such dynamic environments, standard equilibrium concepts such as Markov Perfect Equilibrium (MPE) may admit outcomes in which the exploiter optimally extracts too aggressively, driving the state into collapse despite positive continuation payoffs. These equilibria rely on the exploiter being indifferent to catastrophic breakdown once it occurs. To address this, this paper introduces the Sustainable Exploitation Equilibrium (SEE), a refinement of stationary Markov Perfect Equilibrium that rules out such collapse outcomes by deriving viability from rationality. When exiting a sustainability region entails sufficiently large losses for the exploiter—and when a feasible safe action exists that preserves the relationship—any equilibrium placing positive probability on collapse is sequentially irrational. Under these conditions, restricting attention to viable stationary equilibria is without loss.
SEE then further refines the set of rationally viable stationary MPE by imposing renegotiation-proofness and exploiter-optimal selection. SEE is motivated by settings where one party cannot exit but remains responsive, such as in hegemon–client dynamics, as the exploitee retains a strategic effort margin that shapes both payoffs and dynamics. In this way, SEE formalizes the idea of “sustainable exploitation”: the exploiter extracts as much as possible while ensuring the long-run survival and productivity of the exploitee. This rational-viability foundation distinguishes SEE from equilibrium selection rules that impose sustainability as an external constraint.
@@ 3 Model and Definition
Rational interpretation of viability.
We interpret transitions outside a sustainability set $V\subset S$ as catastrophic breakdown for the exploiter, in the sense that sufficiently small reductions in extraction can strictly improve continuation payoffs by preserving the relationship. Under this interpretation, equilibrium behavior that places positive probability on exiting $V$ is incompatible with sequential rationality when collapse entails sufficiently large losses.
@@
Let $V\subseteq S$ be a nonempty compact viability set that is invariant under admissible state transitions. Let $\mathcal{E}$ denote the set of all stationary Markov Perfect Equilibria, and let $\mathcal{E}_{V}\subseteq\mathcal{E}$ denote the subset whose induced state process remains in $V$. Under rational viability (formalized below), restricting attention to $\mathcal{E}_{V}$ is without loss.
@@ 4 Relation to Nash and Markov Perfect Equilibrium
5 Existence under rational viability
We derive viability from rationality by interpreting exits from a sustainability set as catastrophic for the exploiter. We interpret the catastrophe penalty $M$ as a reduced-form representation of franchise loss, regime collapse, or irreversible destruction of the exploitee following breakdown of the relationship.
Assumption (Catastrophic collapse and safe action). Payoffs are bounded: $|u^{X}(s,x,e)|\leq\bar{u}<\infty$. Let $V\subset S$ be a nonempty compact set. For each $s\in V$, there exists a safe action $\underline{x}(s)\in X(s)$ such that, under the exploitee’s best response,
$Q\!\left(V\mid s,\underline{x}(s),\sigma_{E}(s,\underline{x}(s))\right)=1.$
Theorem 1 (Existence of SEE under rational viability).
Suppose the state space $S$ and action sets $X(s),E(s)$ are compact, payoffs are bounded and continuous, and the transition kernel $Q$ satisfies the Feller property. Interpret $S\setminus V$ as a collapse region generating a catastrophic loss for the exploiter.
Then there exists $M^{\star}<\infty$ such that, for all catastrophe penalties $M\geq M^{\star}$, every stationary Markov Perfect Equilibrium of the penalized game
$\hat{u}^{X}(s,x,e)\equiv u^{X}(s,x,e)-M\cdot\mathbf{1}\{s^{\prime}\notin V\},\qquad s^{\prime}\sim Q(\cdot\mid s,x,e),$
is viable. Moreover, among viable stationary Markov Perfect Equilibria, a nonempty compact subset of renegotiation-proof equilibria exists, and an exploiter-optimal selection exists. Hence a Sustainable Exploitation Equilibrium exists.
@@ 5 Existence under rational viability
Under compactness and continuity, stationary Markov Perfect Equilibria exist. Fix any $s\in V$. If an equilibrium action induces $Q(S\setminus V\mid s,x,e)>0$, the exploiter’s expected one-period payoff is at most $\bar{u}-Mp$ for some $p>0$. Under the safe action $\underline{x}(s)$, the expected penalty is zero and the payoff is at least $\bar{u}$. For $M$ sufficiently large, the safe action strictly dominates any action with positive exit probability. By continuity and compactness, a finite threshold $M^{\star}$ exists. Hence any equilibrium must be viable. Existence of renegotiation-proof and exploiter-optimal selections follows by compactness and continuity. ∎
@@ 6.6 Stationary SEE and regimes
A stationary SEE consists of $(x^{*}(s),e^{*}(s,x))$ and an invariant distribution supported on $V=\{s\geq s_{\min}\}$. Under rational viability, two canonical regimes arise:
@@
Additional Remark on Outside Option
One might alternatively assume that the exploitee $E$ has an outside option $U^{out}_{E}(s)$, in which case viability must be supplemented by an individual rationality (IR) constraint: $W^{E}(s)\geq U^{out}_{E}(s)$ for all $s\in V$. This yields an even tighter refinement, as it restricts feasible equilibria to those in which the exploitee weakly prefers participation over exit.
@@ 7 Conclusion
With a non-exiting but strategically responsive exploitee, SEE refines Markov Perfect Equilibrium by deriving viability from rationality and imposing renegotiation-proof exploiter-optimal selection. Unlike MPE, SEE eliminates equilibria in which the exploiter optimally drives the state to collapse despite positive continuation payoffs. In equilibrium, SEE yields either an interior “golden-rule”–type extraction path or, when unconstrained extraction would cause collapse, a boundary outcome in which the state is pinned at the viability threshold. The framework applies broadly to settings in which one party cannot exit but retains an effort margin—such as colonial relationships, labor arrangements, and locked-in suppliers—providing a compact refinement for analyzing persistent exploitative interactions in economics and political economy.